OpenStudy (anonymous):

Write the expression using factorials: 30 . 29

OpenStudy (anonymous):

\[30\cdot29=\frac{30\cdot29\cdot28}{28}=\frac{30\cdot29\cdot28\cdot27}{28\cdot27}=\cdots=\frac{30!}{28!}\]